A card is drawn at a random from a pack of 52 playing cards (without jokers). Find the probabilities of the
following events happening:
A queen is drawn.

Answer by jitendra_maths(6)

since 4 queens are there in the pack of playing cards.
Hence P(of drawing a queen) = 4/52 =1/13